OpenStudy (anonymous):

can anyone help me??? medal for 1st answer!! x-10=2^2

OpenStudy (anonymous):

so if x-10=2^2 (which is 4), then x=14, because you add 10 to both sides

OpenStudy (hockeychick23):

ok so: x-10=2^2 2^2=4 x-10=4 add 10 to both sides x-10+10=4+10 x=14

OpenStudy (anonymous):